# The InSAR absolute phase amid singularities

### v1

@@ 4.3.3 Intermittent decorrelation
The persistent contribution $u_{\mathrm{p}}$ undergoes an $s$-dependent phase change. The nuisance contribution $u_{\mathrm{n}}$ induces intermittent decorrelation because it is modulated by $f(s)=16s(1-s)$. Both $u_{\mathrm{p}}$ and $u_{\mathrm{n}}$ are drawn independently from a circularly-symmetric standard complex normal distribution and remain perfectly coherent [1].
The absolute phase $\phi^{\mathrm{a}}(s)$ obtained from the ensemble coherence for fixed primary acquisition at $s=0$ is shown by a thick black line in Fig. 4. Due to the persistent contribution, $\phi^{\mathrm{a}}$ increases linearly with $s$, reaching $\pi$ at $s=1$. In the polar plot on the right, the black line starts at $\rho(0)=1$, $\varphi(0)=0$, moving to the left with increasing $s$ as the phase increases while the coherence magnitude $\rho$ first drops before recovering. For $s=1$, the wrapped phase $\varphi(1)=\pi$ would be congruent with an absolute phase of $-\pi$ or $\pi$, but $\phi^{\mathrm{a}}(1)=\pi$ due to the trajectory.
@@
The impact of speckle on the absolute phase is illustrated by the thin lines in Fig. 4, corresponding to multiple realizations of 8-look observations. Most realizations – shown in gray – end at an absolute phase $\phi^{\mathrm{a}}(1)=\pi$, while some – shown in red – terminate at $-\pi$. In the polar plot, the red lines are those that pass the imaginary axis below the origin. Despite the perfect coherence $\rho=1$ and error-free $\varphi=\pi$ observations for $s=1$, the absolute phase differs between realizations due to the intermittent decorrelation. The absolute phase is not amenable to ensemble averaging. If the realizations were from different locations within a homogeneous region, the absolute phase would be discontinuous.
@@ 4.3.4 Tree canopy observations
Transient decreases in coherence are also observed in the L-band BorealScat tomographic time series of a forest canopy in Remningstorp, southern Sweden [24, 25]. Fig. 5a) shows an uncalibrated tomogram from July 24, 2018, derived from the tower-mounted antenna array [26], with the strongest return from the top of the canopy 20–25 m above the forest floor. The upper canopy has $\rho\gtrsim 0.8$ in the unreferenced one-day interferogram in b), from midnight to midnight, and the wrapped phase $\varphi$ is largely within $\pi/8$ of zero.
The estimated absolute phase in Fig. 5c) shows $2\pi$ discontinuities, despite the near uniformity of the 24-hour interferogram. The absolute phase was estimated by approximating the integral (17) by adding the wrapped phase differences over successive 5-minute intervals, the sampling interval of the BorealScat observations. Upper-canopy points A and B differ by $2\pi$ in this estimated absolute phase, although the 24-hour complex coherence is similar. The complex plane insets shows that the coherence timeseries (for fixed primary) differ between these points, with only that of A encircling the origin during a period of low coherence. The low-coherence period occurred during the day, characterized by windier conditions and reduced plant water content. The interpretation is challenging because i) unwrapping the phase at 5-min intervals along the absolute path can introduce errors, and ii) the absolute phase of a speckle-affected realization differs from that of the hypothesized speckle-free underlying process (Fig. 4). Despite these challenges, assuming that the observations sample a coherence that varies smoothly in space and time and that the phase is adequately sampled in time, the absolute phase jumps of $2\pi$ in Fig. 5c) coincide with absolute phase singularities for which the coherence vanished at an earlier, intermediate time.
@@ 5.1 Definition
For monostatic InSAR with a non-zero across-track spatial baseline [3, 2], we will distinguish single-pass and repeat-pass observations in the definition of $\Gamma^{\mathrm{a}}$ through (15). In the single-pass case with $t_{1}=t_{2}$, the primary and secondary configuration only differ in the spatial baseline. For a fixed line-of-sight direction, consider a simplified two-dimensional scenario in which the parallel baseline is zero. Denoting the perpendicular antenna position by $b$, we have $\mathbf{w}=b$ and the perpendicular baseline $b_{2}-b_{1}$. The path $\Gamma^{\mathrm{a}}$ traces antenna positions from $b_{1}$ to $b_{2}$. In the repeat-pass case with $t_{1}\neq t_{2}$, we have $\mathbf{w}=[t,b]$. The natural path $\Gamma^{\mathrm{a}}$ from (15) uniformly and simultaneously increases the temporal and spatial baseline. We can think of it as the antenna moving slowly from primary to secondary position. In both single-pass and repeat-bass cases, a frequency offset $\Delta f$ applied to mitigate surface decorrelation [1] could be incorporated by appending $\Delta f$ to $\mathbf{w}$. Conversely, the following examples focus on single-look observations of point-like scatterers.
@@ 5.2.1 Single-pass: Point scatterer
The single-pass absolute phase of an ideal point scatterer equals the difference in propagation phases [3]. Denoting the distance between the scatterer and the antenna at $b$ by $R(b)$, the single-look-complex signal, normalized to unit magnitude, is $u_{\mathrm{point}}(b)=\exp\left(-\mathsf{j}2kR(b)\right)$. According to (15), the absolute phase is obtained from the series of interferograms with baseline $s\left(b_{2}-b_{1}\right)$, where $s$ is increased from 0 to 1. The integral yields
$\phi^{\mathrm{a}}_{\mathrm{point}}=-\mathsf{j}2k\left(R(b_{1})-R(b_{2})\right)\,.$ (24)
This is just the difference in the propagation phases and is proportional to the range difference $R(b_{1})-R(b_{2})$. It coincides with prior work that stipulated the absolute phase to be equal to this difference [3, 2].
The absolute phase of a point scatterer is essential for height determination [3]. For illustration purposes, we determine the (normalized) view-perpendicular position $q$, which can be converted to a height, and keep the along-axis position $p$ fixed. To simplify notation, let $b_{1}=0$ and interpret $b_{2}\equiv b$ as the spatial baseline, expressed in arbitrary units. Fig. 6a) shows $\phi^{\mathrm{a}}$ of the point scatterer as a function of $b$ for different values of $q$. For any baseline $b$, $\phi^{\mathrm{a}}$ depends on $q$. Estimating $q$ from $\phi^{\mathrm{a}}$ is tantamount to finding $q$ such that the point scatterer’s modeled $\phi^{\mathrm{a}}_{\mathrm{point}}$ equals $\phi^{\mathrm{a}}$. This estimate is commonly interpreted as a phase center position corresponding to the observed phase [2].
Figure 6: Single-pass absolute phase for single-look observations of a) one or b) two point scatterers versus across-track baseline $b$, non-dimensionalized such that the Fresnel number is 1. In a), the lines correspond to across-track positions $q$ (nondimensionalized). In b), the two scatterers are located at $q^{\circ}=0$ and $q^{\bullet}=1$; the heavy lines correspond to different weights $\alpha$.
@@ 5.2.2 Single-pass: Two point scatterers
The two-scatterer absolute phase is plotted as a function of baseline $b$ with heavy lines in Fig. 6b). For equally strong scatterers ($\alpha=0.5$; black), the absolute phase $\phi^{\mathrm{a}}(b)$ for small baselines $b\lesssim 0.6$ agrees with that of a point scatterer located at $q=0.5$, halfway between the two scatterers. However, negative interference in the single-look interferogram for increasing $b$ is seen in the polar plot. For a critical baseline $b_{\mathrm{c}}\approx 0.6$, the circle in both plots demarcates where the return is numerically zero. At this singularity, $\phi^{\mathrm{a}}$ of a point scatterer at $q=0=q_{\circ}$ deviates from that of a point scatter at $q=1=q^{\bullet}$ by $\pi$. While $\varphi$ is well-defined for $b>b_{\mathrm{c}}$, neither the absolute phase nor the phase center position are.
@@ 5.2.3 Repeat-pass: Two point scatterers
Fig. 7 shows a dynamic scenario, in which the scattering strength $\alpha(t)$ depends on time $t$. Starting at $\alpha(0)=0.6$ at the primary time $t_{1}=0$, it dips below 0.6, returning to 0.6 for secondary time $t_{2}=1$. The absolute phase $\phi^{\mathrm{a}}(1)$ at the endpoint of the absolute path $\Gamma^{\mathrm{a}}$ to $\mathbf{w}_{2}=[1,1]^{T}$ is approximately -5 in this dynamic case, while the static case with constant $\alpha=0.6$ is greater by $2\pi$. The difference in absolute phase stems from the observation in the polar plot inset that the $\gamma(t)$ trajectories pass the origin on opposite sides. Conversely, the complex coherence and thus $\varphi$ at the $\Gamma^{\mathrm{a}}$ endpoint are identical for the dynamic and static case. The path $\Gamma^{\mathrm{a}}$ is shown Fig. 7c–d) on top of $\varphi$ and $\rho$ for the dynamic case as a function of variable secondary time $t$ and baseline $b$. Two phase singularities are seen, around which the phase $\varphi$ spirals by $2\pi$.
Figure 7: Repeat-pass InSAR model of two point scatterers with scattering weight $\alpha$. The weights $\alpha(t)$ for the dynamic and a static case (a) correspond to absolute phases (b) that differ by $2\pi$ at $t=1$, as their $\gamma$ pass the origin at opposite sides (inset). The wrapped phase $\varphi$ (c), coherence magnitude $\rho$ (d), and absolute phase $\phi^{\mathrm{a}}$ (e) for variable secondary time $t$ and baseline $b$.
@@ 6 Patching absolute phase singularities
Fig. 7e features absolute phase singularities across which the absolute phase jumps by a non-zero integer multiple of $2\pi$. These absolute phase singularities cannot simply be patched. By patching, I mean extending $\phi^{\mathrm{a}}$ continuously from $W^{\mathrm{a}}$ to $W^{+}$ to obtain a function $\psi:W^{+}\rightarrow\mathbb{R}$ such that $\psi(\mathbf{w})=\phi^{\mathrm{a}}(\mathbf{w})$ for all $\mathbf{w}\in W^{\mathrm{a}}$ and $\mathrm{d}\psi=d\varphi_{W^{+}}$ in $W^{+}$. Loosely speaking, the goal is to add an integer multiple of $2\pi$ to the wrapped phase at absolute phase singularities for which the wrapped phase is defined, $A^{\prime}=A\setminus{\gamma^{-1}\left[\{0\}\right]}$, such that the resulting continuous function $\psi$ matches the $\phi^{\mathrm{a}}$ on $W^{\mathrm{a}}$.
[…]
@@ 6.1 Patching and path independence
[…]
To see the if part, define $\psi(\mathbf{w}_{2})$ as the integral of $d\varphi_{W^{+}}$ along any path in $W^{+}$ from $\mathbf{w}_{1}$ to $\mathbf{w}_{2}$. The function $\psi$ is well defined due to path independence and continuous. For $\mathbf{w}_{2}\in W^{\mathrm{a}}$, $\psi(\mathbf{w}_{2})=\phi^{\mathrm{a}}(\mathbf{w}_{2})$. However, $\psi$ can be evaluated at absolute phase singularities $A^{\prime}$. Basically, for $\mathbf{w}_{2}\in A^{\prime}$, we have replaced the problematic $\Gamma^{\mathrm{a}}$ by another path. It can be shown that $d\varphi_{W^{+}}=\mathrm{d}\psi$ in $W^{+}$, as required [21].
@@ 6.2 Loops for characterizing path dependence
To identify path dependence of the phase integral, the concept of loops introduced in Sec. 2.2.3 is convenient [10]. For arbitrary $\mathbf{w}_{2}$, we consider two paths $\Gamma$ and $\Gamma^{\prime}$, i.e., functions $[0,1]\rightarrow W^{+}$ with $\Gamma(0)=\Gamma^{\prime}(0)=\mathbf{w}_{1}$ and $\Gamma(1)=\Gamma^{\prime}(1)=\mathbf{w}_{2}$. To form a loop, consider the inverse path $\bar{\Gamma^{\prime}}$, defined by $\bar{\Gamma^{\prime}}(s)=\Gamma^{\prime}(1-s)$: the traversal is in the opposite direction from $\mathbf{w}_{2}$ to $\mathbf{w}_{1}$. Concatenation $\Gamma\cdot\bar{\Gamma^{\prime}}$ yields a loop that starts and ends at $\mathbf{w}_{1}$, the loop’s basepoint. This construction [8] is useful because the integral in (12) is path independent for all $\mathbf{w}_{2}$ if and only if
@@
[…]
where $G$ is an oriented surface whose boundary is the loop $g$ with basepoint $\mathbf{w}_{1}$. If we can find such a $G$, the right-hand side vanishes because $\mathrm{d}d\varphi_{W^{+}}=0$, implying path-independence. Deforming $g$ continuously to $\mathbf{w}_{1}$ – essentially pulling $g$ tight like a lasso – implicitly defines such a surface. Such a continuous deformation is called a homotopy [8]. If it exists, we say $g$ is homotopic to $\varepsilon_{\mathbf{w}_{1}}$, the constant loop at $\mathbf{w}_{1}$, with the basepoint $\mathbf{w}_{1}$ held fixed. This is denoted by $g\simeq\varepsilon_{\mathbf{w}_{1}}$.

### v3

@@ 4.3.3 Intermittent decorrelation
The persistent contribution $u_{\mathrm{p}}$ undergoes an $s$-dependent phase change. The nuisance contribution $u_{\mathrm{n}}$ induces intermittent decorrelation: The modulation by f(s)=16s(1s)f(s)=16s(1-s) is chosen to vanish at $s=0$ and $s=1$. Both $u_{\mathrm{p}}$ and $u_{\mathrm{n}}$ are drawn independently from a circularly-symmetric standard complex normal distribution and remain perfectly coherent [bamler98].
The absolute phase $\phi^{\mathrm{a}}(s)$ obtained from the ensemble coherence for fixed primary acquisition at $s=0$ is shown by a thick black line in Fig. 4. Due to the persistent contribution, $\phi^{\mathrm{a}}$ increases linearly with $s$, reaching $\pi$ at $s=1$. In the polar plot on the right, the black line starts at $\rho(0)=1$, $\varphi(0)=0$, moving to the left with increasing $s$ as the phase increases while the coherence magnitude $\rho$ first drops before recovering. For $s=1$, the wrapped phase $\varphi(1)=\pi$ would be congruent with an absolute phase of $\pi$ or $\pi$, but $\phi^{\mathrm{a}}(1)=\pi$ due to the trajectory.
@@
The impact of speckle on the absolute phase is illustrated by the thin lines in Fig. 4, corresponding to multiple realizations of 8-look observations. Most realizations – shown in gray – end at an absolute phase $\phi^{\mathrm{a}}(1)=\pi$, while some – shown in red – terminate at $\pi$. In the polar plot, the red lines are those that pass the imaginary axis below the origin. Despite the perfect coherence $\rho=1$ and error-free $\varphi=\pi$ observations for $s=1$, the absolute phase differs between realizations due to the intermittent decorrelation. The absolute phase is not amenable to ensemble averaging. If the realizations were from different locations within a homogeneous region, the absolute phase would be discontinuous.
@@ 4.3.4 Tree canopy observations
Transient coherence drops are also observed in the L-band BorealScat tomographic time series of a forest canopy in Remningstorp, southern Sweden [monteith22, esa23]. Fig. 5a) shows an uncalibrated tomogram from July 24, 2018, derived from the tower-mounted antenna array [ulander18], with the strongest return from the top of the canopy 20–25 m above the forest floor. The upper canopy has $\rho\gtrsim 0.8$ in the unreferenced one-day interferogram in b) and the wrapped phase $\varphi$ is largely within $\pi/8$ of zero.
The estimated absolute phase in Fig. 5c) shows $2\pi$ discontinuities, despite the near uniformity of the 24-hour interferogram. The absolute phase was estimated by approximating the integral (17) by adding the wrapped phase differences over successive 5-minute intervals, the sampling interval of the BorealScat observations. Upper-canopy points A and B differ by $2\pi$ in this estimated absolute phase, although the 24-hour complex coherence is similar. The complex plane insets shows that the coherence timeseries (for fixed primary) differ between these points, with only that of A encircling the origin during a period of low coherence. The low-coherence period occurred during the day, characterized by windier conditions and lower plant water content.
The interpretation of the estimated absolute phase is challenging. The first challenge is unwrapping errors along the absolute path, despite the short 5-min sampling interval. Second, the absolute phase of a speckle-affected realization differs from that of the hypothetical speckle-free underlying process (Fig. 4). Under the assumption that the observations sample a coherence that varies smoothly in space and time and that the phase is adequately sampled in time, the absolute phase jumps of $2\pi$ in Fig. 5c) coincide with absolute phase singularities for which the coherence vanished at an earlier, intermediate time. The jumps reflect the sensitivity of phase unwrapping to perturbations at low coherence, and their occurrence in integer multiples of $2\pi$ is due to the congruence of the absolute phase with the (near-zero) wrapped phase. Both characteristics impede interpreting the observed absolute phase as a range difference.
@@ 5.1 Definition
For monostatic InSAR with a non-zero across-track spatial baseline [madsen93, rosen00], we will distinguish single-pass and repeat-pass observations in the definition of $\Gamma^{\mathrm{a}}$ through (15). In the single-pass case with $t_{1}=t_{2}$, the primary and secondary configuration only differ in the spatial baseline. For a fixed line-of-sight direction, consider a simplified two-dimensional scenario in which the parallel baseline is zero. Denoting the perpendicular antenna position by $b$, we have $\mathbf{w}=b$ and the perpendicular baseline $b_{2}-b_{1}$. The path $\Gamma^{\mathrm{a}}$ traces antenna positions from $b_{1}$ to $b_{2}$. In the repeat-pass case with $t_{1}\neq t_{2}$, we have $\mathbf{w}=[t,b]$. The natural path $\Gamma^{\mathrm{a}}$ from (15) uniformly and simultaneously increases the temporal and spatial baseline. We can think of it as the antenna moving slowly from primary to secondary position. In both single-pass and repeat-pass cases, a frequency offset $\Delta f$ applied to mitigate surface decorrelation [bamler98] could be incorporated by appending $\Delta f$ to $\mathbf{w}$. Conversely, the following examples focus on single-look observations of point-like scatterers.
@@ 5.2.1 Single-pass: Point scatterer
The single-pass absolute phase of an ideal point scatterer equals the difference in propagation phases [madsen93]. Denoting the distance between the scatterer and the antenna at $b$ by $R(b)$, the single-look-complex signal, normalized to unit magnitude, is $u_{\mathrm{point}}(b)=\exp\left(-\mathsf{j}2kR(b)\right)$. According to (15), the absolute phase is obtained from the series of interferograms with baseline $s\left(b_{2}-b_{1}\right)$, where $s$ is increased from 0 to 1. The integral yields
$\phi^{\mathrm{a}}_{\mathrm{point}}=-2k\left(R(b_{1})-R(b_{2})\right)\,.$ (24)
This is just the difference in the propagation phases and is proportional to the range difference $R(b_{1})-R(b_{2})$. It coincides with prior work that stipulated the absolute phase to be equal to this difference [madsen93, rosen00].
The absolute phase of a point scatterer is essential for height determination [madsen93]. For illustration purposes, we determine the view-perpendicular position $q$, which can be converted to a height, and keep the along-axis position $p$ fixed. To simplify notation, let $q$ be normalized (with $q=1$ corresponding to a target Fresnel number $F_{q}=1$), let $b_{1}=0$ and interpret $b_{2}\equiv b$ as the spatial baseline, expressed in arbitrary units. Fig. 6a) shows $\phi^{\mathrm{a}}$ as a function of $b$ for different values of $q$. For any baseline $b$, $\phi^{\mathrm{a}}$ depends on $q$. Conversely, an estimate of $q$ is obtained by matching the observed $\phi^{\mathrm{a}}$ to the point scatterer’s modeled $\phi^{\mathrm{a}}_{\mathrm{point}}$. This estimate is commonly interpreted as a phase center position corresponding to the observed phase [rosen00].
Figure 6: Single-pass absolute phase for single-look observations of a) one or b) two point scatterers versus across-track baseline $b$, non-dimensionalized such that the basline Fresnel number $F_{b}=0.16$. In a), the lines correspond to across-track positions $q$ (nondimensionalized). In b), the two scatterers are located at $q^{\circ}=0$ and $q^{\bullet}=1$; the heavy lines correspond to different weights $\alpha$.
@@ 5.2.2 Single-pass: Two point scatterers
The two-scatterer absolute phase is plotted as a function of baseline $b$ with heavy lines in Fig. 6b). For equally strong scatterers ($\alpha=0.5$; black), the absolute phase $\phi^{\mathrm{a}}(b)$ for small baselines $b\lesssim 0.6$ agrees with that of a point scatterer located at $q=0.5$, halfway between the two scatterers. However, the inset plots shows decreasing magnitude with increasing $b$ (direction of arrows) arising from negative interference, while the magnitude of a single scatterer would be independent of baseline. For a singular baseline $b_{\mathrm{s}}\approx 0.6$, the circle in both plots demarcates where the return is numerically zero. At this singularity, $\phi^{\mathrm{a}}$ of a point scatterer at $q=0=q_{\circ}$ deviates from that of a point scatter at $q=1=q^{\bullet}$ by $\pi$. While $\varphi$ is well-defined for $b>b_{\mathrm{s}}$, neither the absolute phase nor the phase center position are.
@@ 5.2.3 Repeat-pass: Two point scatterers
Fig. 7a–b shows a dynamic scenario, in which the scattering strength $\alpha(t)$ depends on time $t$. Starting at $\alpha(0)=0.6$ at the primary time $t_{1}=0$, it dips below 0.6, returning to 0.6 for secondary time $t_{2}=1$. The absolute phase $\phi^{\mathrm{a}}(1)$ at the endpoint of the absolute path $\Gamma^{\mathrm{a}}$ to $\mathbf{w}_{2}=[1,1]$ is approximately -5 in this dynamic case, while the static case with constant $\alpha=0.6$ is greater by $2\pi$. The difference in absolute phase stems from the observation in the polar plot inset that the coherence trajectories along the absolute path pass the origin on opposite sides. Conversely, the complex coherence and thus $\varphi$ at the $\Gamma^{\mathrm{a}}$ endpoint are identical for the dynamic and static case. Fig. 7c–d) show $\varphi$ and $\rho$ for the dynamic case as a function of variable secondary time $t$ and baseline $b$. The absolute path, which is independent of the observed target, passes between two phase singularities, around which the phase $\varphi$ spirals by $2\pi$.
Figure 7: Repeat-pass InSAR model of two point scatterers with scattering weight $\alpha$. The weights $\alpha(t)$ for the dynamic and a static case (a) correspond to absolute phases (b) that differ by $2\pi$ at $t=1$, as their $\gamma$ pass the origin at opposite sides (inset). The wrapped phase $\varphi$ (c), coherence magnitude $\rho$ (d), and absolute phase $\phi^{\mathrm{a}}$ (e) for variable secondary time $t$ and baseline $b$, and dynamic $\alpha(t)$.
@@ 6 Patching absolute phase singularities
Fig. 7e features absolute phase singularities across which the absolute phase jumps by a non-zero integer multiple of $2\pi$. These cannot be patched. By patching, I mean extending $\phi^{\mathrm{a}}$ continuously from $W^{\mathrm{a}}$ to $W^{+}$ to obtain a function $\psi:W^{+}\rightarrow\mathbb{R}$ such that $\psi(\mathbf{w})=\phi^{\mathrm{a}}(\mathbf{w})$ for all $\mathbf{w}\in W^{\mathrm{a}}$ and $\mathrm{d}\psi=d\varphi_{W^{+}}$ in $W^{+}$. Loosely speaking, the goal is to add an integer multiple of $2\pi$ to the wrapped phase at absolute phase singularities such that the resulting continuous function $\psi$ matches $\phi^{\mathrm{a}}$ off the absolute phase singularities. More precisely, we can only hope to patch those absolute phase singularities for which the wrapped phase is defined, i.e., $A^{\prime}=A\setminus{\gamma^{-1}\left[\{0\}\right]}$.
An example of a patchable absolute phase singularity is shown in Fig. 8a. The absolute phase $\phi^{\mathrm{a}}$ is the same on either side of the absolute phase singularity $A$ for fixed $w_{2,2}$, allowing us to extend $\phi^{\mathrm{a}}$ continuously. Conversely, the absolute phase singularities in Fig. 8b cannot be patched.
Figure 8: Two examples of absolute phase singularities for varying $\mathbf{w}_{2}=[w_{2,1},w_{2,2}]$ with fixed $\mathbf{w}_{1}=[0,0]$ for: a) patchable and b) unpatchable (same example as Fig. 7). The corresponding wrapped phases $\varphi$ are shown in c) and d), respectively, and the loops projected into the complex plane in e) and f).
When can we patch the absolute phase singularities? Patchability turns out to be closely connected to path independence of line integrals $\int d\varphi_{W^{+}}$.
@@ 6.1 Patching and path independence
The absolute phase can be extended across absolute phase singularities $A^{\prime}$ if and only if the integral $\int_{\Gamma}d\varphi_{W^{+}}$ does not depend on the path $\Gamma\in W^{+}\subset\mathbb{R}^{n}$ from fixed $\mathbf{w}_{1}$ to all $\mathbf{w}_{2}$. It is assumed that $\gamma(\mathbf{w}_{2})$ is smooth and $W^{+}\cong S\subset\mathbb{R}^{n}$ is path connected and locally path connected. The statement is closely related to the fact that a closed one-form $\omega$ on a subset of $\mathbb{R}^{n}$ is exact – i.e., there exists a function $f$ such that $\omega=\mathrm{d}f$ – if and only if the line integral $\int_{\Gamma}\omega$ is path independent (Theorem 6.4.12 in Chapter IV, [weintraub14]).
To see the if part, define $\psi(\mathbf{w}_{2})$ as the integral of $d\varphi_{W^{+}}$ along any path in $W^{+}$ from $\mathbf{w}_{1}$ to $\mathbf{w}_{2}$. The function $\psi$ is well defined due to path independence and continuous. For $\mathbf{w}_{2}\in W^{\mathrm{a}}$, $\psi(\mathbf{w}_{2})=\phi^{\mathrm{a}}(\mathbf{w}_{2})$. However, $\psi$ can be evaluated at absolute phase singularities $A^{\prime}$. Basically, for $\mathbf{w}_{2}\in A^{\prime}$, we have replaced the problematic $\Gamma^{\mathrm{a}}$ by another path. It can be shown that $d\varphi_{W^{+}}=\mathrm{d}\psi$ in $W^{+}$, as required [weintraub14].
@@ 6.2 Loops for characterizing path dependence
To identify path dependence of the phase integral, the concept of loops introduced in Sec. 2.2.3 is convenient [hatcher01]. For arbitrary $\mathbf{w}_{2}$, we consider two paths $\Gamma$ and $\Gamma^{\prime}$, i.e., functions $0,1]\rightarrow W^{+}$ with $\Gamma(0)=\Gamma^{\prime}(0)=\mathbf{w}_{1}$ and $\Gamma(1)=\Gamma^{\prime}(1)=\mathbf{w}_{2}$. To form a loop, consider the inverse path $\bar{\Gamma^{\prime}}$, defined by $\bar{\Gamma^{\prime}}(s)=\Gamma^{\prime}(1-s)$: the traversal is in the opposite direction, from $\mathbf{w}_{2}$ to $\mathbf{w}_{1}$. Concatenation $\Gamma\cdot\bar{\Gamma^{\prime}}$ yields a loop that starts and ends at $\mathbf{w}_{1}$, the loop’s basepoint. This construction [fulton97] is useful because the integral in (12) is path independent for all $\mathbf{w}_{2}$ if and only if
@@
In the example of the patchable phase singularity in Fig. 8, consider the loops $g$ and $g^{\prime}$ shown in panel b. Integrating $d\varphi_{W^{+}}$ over $g$ or $g^{\prime}$ yields zero, since the wrapped phase first increases and then decreases along either loop. Indeed, the loop integral vanishes for any loop in $W^{+}$, which is equivalent to patchability. Conversely, for the unpatchable absolute phase singularity, panel d shows a loop $g$ with $\int_{g^{\prime}}d\varphi_{W^{+}}\neq 0$. Even though the integral vanishes over the other loop $g^{\prime}$, the existence of a single loop with non-zero integral implies the absolute phase cannot be patched.
For a loop $g$ in $W^{+}$, $\int_{g}d\varphi_{W^{+}}=0$ if the loop can be continuously deformed to a single point in $W^{+}$. The loops $g^{\prime}$ in Fig. 8b,d are such contractible loops. The integral vanishes because $d\varphi_{W^{+}}$ will be shown to be closed, i.e., its exterior derivative $\mathrm{d}d\varphi_{W^{+}}=0$. Applying the exterior derivative $\mathrm{d}$ to a one-form generalizes the notion of taking a curl, as can be seen in Stokes’ theorem [weintraub14]
$\int_{g}d\varphi_{W^{+}}=\int_{G}\mathrm{d}d\varphi_{W^{+}}\,$, (26)
where $G$ is an oriented surface whose boundary is the loop $g$ with basepoint $\mathbf{w}_{1}$. If we can find such a $G$, the right-hand side vanishes because $\mathrm{d}d\varphi_{W^{+}}=0$, implying path-independence. Deforming $g$ continuously to $\mathbf{w}_{1}$ – essentially pulling $g$ tight like a lasso – implicitly defines such a surface. Such a continuous deformation is called a homotopy [fulton97]. If it exists, we say $g$ is homotopic to $\varepsilon_{\mathbf{w}_{1}}$, the constant loop at $\mathbf{w}_{1}$, with the basepoint $\mathbf{w}_{1}$ held fixed. This is denoted by $g\simeq\varepsilon_{\mathbf{w}_{1}}$.